OpenStudy (anonymous):

simplify x^2-1/x^2+2x+1

OpenStudy (anonymous):

(x-1)/(x+1)

OpenStudy (anonymous):

\[\frac{ (a-b)(a+b) }{ (a+b)^2}\]

OpenStudy (anonymous):

\[\frac{ (x-1)(x+1) }{ (x+1)(x+1) }\]

OpenStudy (anonymous):

Cancelling out (x+1) in the numerator and the denominator, your simplified expression is: \[\frac{ (x-1) }{ (x+1) }\]